A horizontal vinyl record of mass 0.0856 kg and radius 0.0898 m rotates freely about a vertical axis through its center with an angular speed of 5.82 rad/s and a rotational inertia of 5.42 x 10-4 kg·m2. Putty of mass 0.0267 kg drops vertically onto the record from above and sticks to the edge of the record.What is the angular speed of the record immediately afterwards?Chapter 11, Problem 055 Your answer is partially correct. A horizontal vinyl record of mass 0.0856 kg and radius 0.0898 m rot
Answer #1

answer) by applying conservation of angular moment we have

initial momentum= final momentum

I\omegai=(I+I1) \omegaf

I1= inertia of putty=mr2

so we have

I\omegai=(I+mr2)\omegaf

so \omegaf=I\omegai/I+mr2

now putting the values we get

\omegaf=5.42*10-4*5.82/5.42*10-4+(0.0267*0.08982)=4.17 rad/s

so the answer is 4.17 rad/s or 4.2 rad/s or 4.165 rad/s